Notes from the meeting respectfully submitted.

Sci IPT 17 March 2010

Folks in Garching--quite a few. Alsacia Hills, Peck, Wootten, Nyman, Wilson, Hales, Lucas, Nikolic OSF Zwaan Kartik Sheth K. Morita OSF M. Wright Darrel Emerson Roll not taken after start, others may have joined.

Richard--Feb was a poor weather month, lots of water got into things. Still struggling on software fronts, particularly with the correlator. A way has been found around these though it was startling to hit software limits at this early stage. Please see the pictures, ALMA Status ASAC. Cost to complete--was thought to be a constraining impetus but currently there is a reasonable amount of contingency and the project can be completed to cost. For ops, there are very big pressure, owing to the currency fluctuations, underestimates on power and road maintenance and other factors. Some costs, exchange rate changes for example, and fuel costs, could be taken into consideration by the funding agencies. Reconstruction will also have an impact. Leaving out intermediate configurations during the expansion and contraction of the array is being considered as one cost savings measure. See readying of the ACA for antennas, the power system will be tested today. Move is scheduled on Saturday. Power is generated locally now. The 23 kV system is complex and needs considerable assurances before it can be started. Turbines projected for next year. Operations are run from the control room, a very warm environment, and access the archive. Four antennas are now at the OSF, two have receivers a third gets one next week, two are run as interferometer. PM02 surface was set by NAOJ to 7.2mm on its surface. At the OSF it will be reset with the appropriate thermal and elevation bias. First 7m antenna is being tested and operated. First of AEM antennas is now assembled, still being outfitted with drives. This has kept to schedule for the last month or so. The surface is put on before the lift and was good to 30microns then. Nest slide shows deliveries. Many items are in good supply. Latest items are cal devices and FEs. Some FE acceleration has occurred, next three to be here by end of May, leaving cal devices as a pacing item. Predictions show supply should be good by end of construction. Band 7 - recall high xpol on original design which was changed after preproduction receivers to a design with a rotated waveguide. The plan was that eight original receivers would get us through initial CSV. But we now have one of those on an antenna. The new guideline is to use the new design when possible and refurbish the older ones when possible. With the mixed mode still get 80% coherence. Still OPT problems--thought that the thermal instability occurs in the aluminum lens mount. The one we have has a burnt out motor. Nutators are moving slowly and ASIAA has taken over servo system. John Ford of GB has assisted in the design but this is still lagging. An alternative is to fast scan the whole dish; we are working on that now. Rx instability though requires moving across the source in 0.1s then turnaround which limits sensitivity. A more gentle turnaround could be implemented--lissajous patterns for example. Also could modify nutator chop time, now at 10ms rise time. Move on to commissioning progress slides. Pointing. A deformation in the base owing to its setting on its kinematic mount (slide 5) caused the problems seen in slide 4. Picking the antenna up and resetting resulted in improvement seen in plot 6. An improvement in procedure has been implemented. Mel says they had to do a similar thing with BIMA antennas. PM03 also had a different type of error having to do with the metrology system which was being used. Cure implemented. Tracking errors seen in slide 10 shows some tracking errors. This has also been done interferometrically. Astronomical holography is working but the results are not meaningful as atmosphere dominates on long baselines. Scans on sun and moon illustrate small scale surface errors which cause the sidelobes and are not measured by holography. Diffraction from legs dominates at low frequencies. Can do this now at higher frequencies. Line length correction system basically works but there are features seen. We are checking the phase at the antenna to check. Phase glitches showed up and then vanished (slide 18) and may have been software but since it is spectra in slide 20 with recent slide 21. Slide 22 shows a line scan in Orion B3 but some normalization problems are evident as the two polzns blue and red should lie atop each other. Full polzn data taken on a quasar see plot 24. One can see the rotation of the combination of source and instrumental polzn. There is a curious effect in the YY polzn, being investigated.

Change requests. Please see Technical Issues presentation. Some details of antenna performance from an analysis by Emerson Schwab and others illustrates some differences in the antennas associated with the subreflector support system. The finite element analysis did not predict these but the surface should perform to spec even at the most extreme conditions. Should achieve 15-20 mucrons normally. Have not measured gravitational terms. Vertex slide 17 shows some settling taking a long time to settle. Outside offset pointing spec but this is after a big slew. Solution may be to make sure inclinometers settle better through reworking or replacement. Could make sure antenna is settled before critical pointing measurements. Front End slide 19. We have CRE-230: Change to IF Bandwidth for Band 6 Cartridges which covers cold cartridges noise and gain slope but, as discussed last time does not include the excess noise on the LO. Need to characterize effect--spurious features or just excess noise? For simultaneous 12CO and 13CO one might want to move between the two lines in a set cadence. – Also CRE-244: Change to Cross Polarization Isolation Specification for Band 6 Cartridges which lowers requirement on polarization purity to 99% (-20dB) from 219 GHz to 230.5 GHz and 98% (-17 dB) at 262 GHz. No limit elsewhere in the band. Spec from FE was -23dB so this departed already. Crystal and Todd looked at various lines and judged it acceptable noting that CCH at 262 GHz and CO/CN at 230 are important. What are the IPT's thoughts on this? They say this will make a 75% yield. Ten delivered, within this range. Some changes in windows and OMT had been made to get to this stage. Worry: we would like some guarantee that there was a best effort being maintained here. Wright--important to get on the air and mess around with them later for improvements. Testi: what is situation with other bands? REH: B 7 is the only one meeting -23dB B3 is around this deteriorating at top end. B9 dominated by beam squint. B4 and B8 are quite close. Probably will see further CREs. REH: I feel that we do need a requirement of -17dB across the whole band. RL: Are these static problems? REH: Now that we have tilted windows standing waves between windows should be eliminated. OMT is also more stable as function of frequency now. If you could look at the CRE Robert, particularly the detailed report, it would be useful. RL: There should be tests of the best version. REH: FEIC scanners had high polzn so this only came to bite us now. B4 CRE: the actual request is not final. Trx 44-60K warm optics 3.8K so the spec should be raised. The request from 51 to 61K is somewhat depressing. IF amps, from commercial rather than Yebes or JPL, not ideal. For B3 the important spec was on overall performance so one can trade off mixer performance to get average which meets spec. Is there a part of the band where we need best performance? Koh-Ichiro Morita: ALMA-J management is worried about this and had meeting yesterday to discuss options. Some further development should come soon. We should await the final change request. REH: Yes we await the final change request. Gain compression is also higher than the 5% spec, could seek to raise this to 9%. Fixed compression can be dealt with but variable compression is difficult. One could trade Trx performance for compression. Probably favor the former. DTE: How reproducible? REH: Pretty steady in the data shown though at the top end 160 GHz it was not, probably associated with Josephson noise. When step is comparable to the Josephson step size there is a problem. Magnetic suppression could be a major redesign. Slide 21 shows a list of CREs in system. Quartz window loss is not a problem. B7 cartridge is a practical issue. B3 one was never approved though discussed a year ago. CRE 234 is a clarification. CRE 244 is the B6 polzn change. Please, Darrel and Robert, examine these and advise.

Procurement of a temperature profiler for ALMA based on radiometric sounding of the O2 lines under way. Contract placement being pushed.

WVRs--next step is the move to 1km baselines which will provide a challenging conditions.

VLTI-ALMA workshop went very well. Report in messenger in June. Also a report on the disk workshop and data needs for ALMA.

April 21st will be the next meeting.

-- AlWootten - 2010-03-17
Topic revision: r1 - 2010-03-17, AlWootten
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback